# Swarm-intelligent search for gravitational waves from eccentric binary mergers

### v3

@@ I Introduction
Nevertheless, significant developments towards eccentric parameter estimation have been made, for example using a hybrid approach- by doing an initial analysis with a non-precessing waveform like the IMRPhenomD and then reweighing the results with the eccentricity-marginalised Bayesian likelihood using the SEOBNRE model (Romero-Shaw et al., 2019, 2020a, 2023, 2021, 2022, 2020b). Studies have indicated that effects due to orbital eccentricity and spin-induced orbital precession can be correlated (Romero-Shaw et al., 2020a, 2023; Bustillo et al., 2021). The lack of accurate and fast-to-generate eccentric waveforms describing the full set of binary parameters hinders their implementation in the LVK analyses. Developing eccentric waveforms with spins is an active area of research in the community. These waveforms are mostly intended for parameter estimation studies. Note that currently, the description of eccentricity is not standardized and can vary across different models (Knee et al., 2022; Shaikh et al., 2023). However, defining eccentricity at 10 Hz is almost a standard practice.
@@
An earlier search targeting non-spinning eccentric BNSs using a stochastic template placement algorithm could not lead to any interesting eccentric BNS candidate (Nitz et al., 2020). If eccentricity is associated with some of the detectable events, a generic eccentric search can enhance the significance of some marginal events flagged by the aligned-spin searches. After this paper was originally submitted, a search appeared for neutron star binaries with a template bank that is about two orders of magnitude larger than the aligned-spin bank (Dhurkunde and Nitz, 2023). That search reports no new significant candidate from the third Observing Run of the Advanced LIGO and Advanced Virgo.
In this work, we demonstrate a spin-aligned eccentric search using the Particle Swarm Optimization (PSO) algorithm. Previously, PSO based searches have been explored towards conducting aligned-spin and even fully precessing searches but never included eccentricity (Wang and Mohanty, 2010; Srivastava et al., 2018; Normandin and Mohanty, 2020). Here, we include the orbital eccentricity as a search parameter in addition to the component masses and their aligned-spins. To generate templates waveforms for BNS systems, we use the inspiral-only TaylorF2Ecc model. It is a fast, frequency-domain model supporting small eccentricities in the presence of aligned-spins (Moore et al., 2016). For injecting signals from simulated sources, we primarily use the eccentric version of the time-domain TEOBResumS model (called TEOBResumS-DALI) that can incorporate very high eccentricities in the presence of aligned-spins (Chiaramello and Nagar, 2020; Nagar et al., 2021). Here we also use it for generating templates for BBH systems. Currently, neither of these waveform models incorporate the mean anomaly parameter and thus, all our injections and the results are obtained at a fixed zero mean anomaly. Our key observations and insights are summarized as follows.
@@
To explore a given search parameter space, the PSO algorithm iteratively computes the templates points on the fly while optimizing a given detection statistic. This approach is different from that of a template-bank which depends on a fixed, pre-determined set of template points. The PSO algorithm thus, avoids the minimal-match criterion implemented in the template-bank searches but spends more time in exploring the region of the parameter space where the signal is more likely to be found. Some parts of the search parameter space generally require larger sampling than others. For example, the signal recovery in the low mass BNS region need more template points than that in the heavier BBHs. When data from more than one detector is available, the PSO algorithm can optimize a statistic computed over the detector network.
@@
Matched-filtering of the data is preceded by template waveform generation in the PSO framework. Eccentric templates are in general costlier than the quasi-circular ones impacting the overall cost of the search. We find that the above strategy picks the aligned-spin and the eccentricity independently from each other. However, we are currently limited by the parameter space coverage of the waveform models and the speed of the dynamic template generation.
@@
Eccentricity primarily modulates the inspiral component of the signals and the gain is thus, from the low frequency content of the signals. On the other end, truncating signals till Hz by downsampling reduces the cost of matched-filtering with little to no impact on recovered the SNRs. If eccentricity is absent in the signals, the eccentric search approximately behaves like the usual aligned-spin search in terms of signal recovery.
@@ II.1 Detection principle
In this section, we briefly review the overall method for coincident detection of compact binary coalescences. The data recorded by a gravitational wave detector is a timeseries of discretely sampled strain, containing occasional gravitational signals in the detector’s noise. The detection problem is to essentially identify instances of transient astrophysical signals in the timeseries data. This is achieved by cross-correlating the strain data with a modeled form of the astrophysical signal, called a template waveform, . A template waveform depends on several parameters, modelling the coalescing binary. The cross-correlation is performed in the frequency domain where the noise power spectral density, is used to down-weight the effect of noise, in a process called matched-filtering. The output of matched-filtering is the signal-to-noise ratio, defined as,
where a tilde represents a function in the frequency domain. Whenever there is a close match between the true signal in the data and a template waveform, the SNR timeseries attains a peak at the arrival time of the signal in the detector as shown in the Fig. 1. However at times, detector noise can also mimic GW signals by causing sharp peaks in the SNR timeseries. These occur mainly due to the non-stationarity and non-Gaussian nature of the detector noise and are called glitches. Glitches increase the chances of false alarms and affect the sensitivity to true astrophysical signals. The peaks in the SNR timeseries above a chosen threshold which arise due to the noise glitches and any potential astrophysical signals are considered as triggers. These undergo several consistency checks to veto out possible glitches. To get rid of some of the obvious loud noise transients, the signal-based power discriminator is used (Allen, 2005). Other vetoing techniques are also available for the noise triggers that dodge the standard consistency checks, especially from short duration templates (Nitz, 2018). The discriminators are designed to down-rank noise triggers in a quantity called the reweighted SNR while preserving the original SNR values for the triggers with astrophysical origin (Usman et al., 2016). Remaining triggers are tested for time-coincidence with multiple detectors - the arrival times of astrophysical signals at any pair of detectors cannot differ by more than the time-of-flight for light between the detectors with some tolerance for the measurement errors. Events that qualify these tests are assigned a statistical significance based on the rate of false alarms, which is described in section II.4. The general detection problem is extensively discussed in the literature, for example (Sathyaprakash and Schutz, 2009; Schutz, 1997; Sathyaprakash and Dhurandhar, 1991).
@@ II.2 Swarm-intelligence
Identifying a potential signal in a given segment of data hinges on finding a close match between the signal and a template waveform. As described in the previous section, one needs to optimize function as given in the Eq. (1) over the desired range of parameters . In this section, we describe how the proposed algorithm locates the optimal template position which is represented by . The algorithm begins with a fixed number of randomly distributed template points which we refer to as a swarm. The individual members of the swarm, which are known as particles, communicate and share the knowledge gained during the search process. Each particle with a position and a velocity , where is the particle identifier and is the dimension index, constitutes a candidate for the optimal template . The initial random distribution of the particles can be chosen to be uniform in each of the search dimensions. The role of the algorithm is to iteratively improve a given fitness function by proposing a new set of positions to the particles and computing the fitness function at the newly proposed locations. The swarm eventually converges over few tens of iterations (typically less than 20) while the value of the fitness function saturates at an optimal value. Note that the fitness function must be carefully chosen since its repeated evaluation can be costly. For example, the simplest choice is the matched-filter SNR maximized over time for a single detector with Gaussian noise. However, a noise glitch can cause a very large SNR value and can deceit the algorithm. Thus we replace the SNR with the reweighted SNR for the network of detectors for maximization, as discussed in Section-II.3. To take any subsequent step, a particle considers the following three factors-
@@
So at any given step, the algorithm identifies two quantities: (a) the template point with the largest fitness function among all the particles, which is called global-best or g-best, denoted as , and (b) the template point with the highest fitness achieved by the particle till the current step, which is called personal-best or p-best, denoted as . The overall velocity at step can be expressed as:
where and are uniformly distributed random numbers between and . The constant weights and control the degree of convergence and exploration of the swarm. A rough set of values for the constants is (, , ) = (0.7, 0.35, 0.4), but fine tuning is possible depending upon the search. The positions of the particles are updated using the simple rule given below,
As an initial test of robustness of the PSO algorithm, we demonstrate the recovery of simulated BNS sources from a small population of about 500 injections in Gaussian noise. These are non-eccentric BNS injections with moderate aligned-spin components; the component masses are drawn from . To locate the optimal template parameters for the HL detector network, we used 5k particles and allowed up to 15 steps, aggregated to about 75k template evaluations and that many matched-filter computations per detector. This recovery corresponds to an aligned-spin search using PSO. A summary of simulation results are shown in the Fig. 2. The injected coincident SNR is the coincident SNR obtained from the template with the injected parameters. Here an average loss of about 0.4% is observed in the recovered coincident SNR from that of the injected values. We also observed that to tackle less dense regions of the search space, like that of the NSBH and BBHs, requires even lesser sampling of the parameter space in general. The final g-best location in the parameter space represents an optimal solution with the obtained template parameters . Here the point estimates of the parameters describing the simulated sources are reasonably unbiased, notably the chirp mass, and the effective spin parameter, as shown in Fig. 2. The modifications to this algorithm to adapt to the real instrumental noise are described in the next section.
@@ II.3 Multiple detector optimization
When searching in the coincident mode, the processes described earlier in section II.2 can, in principle, proceed independently and in parallel for all the available detectors. However, in such a case, the templates are also computed independently for each detector and the point of convergence can differ due to several reasons. When projected from arbirary directions in the sky, the antenna pattern functions of the detetors modulate the incoming signal differently, thus effectively shifting the expected optimal point. In addition, the noise contributions from the individual detectors can slightly shift the convergence point. Otherwise, several local peaks form an elongated band of almost degenerate templates with similar SNRs as shown in the Fig. 3, which can confuse the convergence. But a true gravitational wave signal should have the times of arrival and the source parameters consistent across all the available detectors. To ensure the second criterion, the same set of template parameters should optimize the given fitness function. Besides, for a given template, a stretch of data can contain several noise triggers, however, such triggers are expected to be uncorrelated and appear at just one detector. Thus, to avoid spurious optimization of glitch triggers, a natural choice is to jointly optimize the reweighted SNRs across the network of detectors. The joint fitness function for a network of detectors is given by,
@@
where is the reweighted SNR of the most likely trigger arising from an astrophysical signal at the th detector. Template points that trigger to a loud glitch in one of the detectors are downranked by the veto processes and the swarm further inhibits convergence to such template points over the detector network. However, at this stage, we do not conduct a time coincidence test of the triggers and is performed after the optimization. An additional advantage of the above procedure is that a single set of computation of the template waveforms serves the entire network of detectors. It is especially useful when generating the template waveforms is expensive as is generally the case in an eccentric search. A similar strategy is also employed to optimize the network SNR of the candidates found in the PyCBC Live search (Canton et al., 2021).
@@
To evaluate the fitness function at any given iteration, the template waveform at is constructed and cross-correlated with the given stretch of data. Each of these results into an SNR timeseries whose ends can be corrupted due to filtering artifacts and are discarded. Triggers from the remaining portion of the SNR timeseries above a chosen threshold, typically , are examined for all the particles. The reweighted SNR values are computed at the trigger points using the power -test to reweigh the triggers. The reweighted SNR is designed to downrank the noise triggers while preserving the SNR values of signals modelled by templates. Additionally, we compute the ratio of the reweighted-SNR and the SNR for the triggers and use it to rank them for their astrophysical origin. Triggers with a ratio close to 1, typically above are retained and only the highest ranked trigger from each detector proceed for joint optimization over the network as given in (Eq. 4). Note that these primary triggers do not undergo any time coincidence test and are only used to arrive at the best guess of the g-best template over the detector network. Thus we proceed with the optimization to get the final g-best parameters and store the resulting triggers indefinitely. The coincidences are formed only using these secondary triggers from the final g-best template in a post-processing step. This is in contrast to the template-bank searches where each template in the bank can give rise to a foreground candidate and hence, also participate in the background estimation. However, it is still consistent that the triggers which are tested for foreground coincidences are used to estimate the background.
@@
After updating the position-velocity, the algorithm also checks whether any particle has drifted beyond the desired search boundaries- a reflecting boundary puts back an outlier inside in a symmetric position while flipping its velocity. All particles undergo this condition after the newly proposed template positions using Eq. (3) is computed. While this is suitable in most of the cases, we find that an adsorbing, boundary is more effective for the eccentricity dimension towards the recovery of non-eccentric systems where instead of putting the outlier back to a symmetric location, we put it on the boundary. At this point, the current iteration is complete. The algorithm additionally checks for a terminal condition such as a convergence below a small mean radius of the swarm or a preset timeout. The boundary conditions can improve the termination and also ensures no hang-up occurs due to non-availability of template parameters outside the region allowed by a waveform model. Note that these steps, namely evolving the position-velocity, applying the boundary conditions and the termination checks are computationally trivial. A summary of the steps are shown in the flow chart in Fig. 4.

### v5

@@ I Introduction
An earlier search targeting non-spinning eccentric BNSs using a stochastic template placement algorithm could not lead to any interesting eccentric BNS candidate (Nitz et al., 2020). If eccentricity is associated with some of the detectable events, a generic eccentric search can enhance the significance of some marginal events flagged by the aligned-spin searches. After this paper describing spin-aligned eccentric searches was originally submitted, a similar search appeared for neutron star binaries but with a template bank that is about two orders of magnitude larger than a corresponding aligned-spin bank (Dhurkunde and Nitz, 2023). That search reports no new significant candidate from the third Observing Run of the Advanced LIGO and Advanced Virgo.
In this work, we demonstrate a spin-aligned eccentric search using the Particle Swarm Optimization (PSO) algorithm. Previously, PSO based searches have been explored towards conducting aligned-spin and even fully precessing searches but never included eccentricity (Wang and Mohanty, 2010; Srivastava et al., 2018; Normandin and Mohanty, 2020). Here, we include the orbital eccentricity () as a search parameter in addition to the component masses and their aligned-spins. To generate template waveforms for BNS systems, we use the inspiral-only TaylorF2Ecc model. It is a fast, frequency-domain model supporting small eccentricities in the presence of aligned-spins (Moore et al., 2016). For injecting signals from simulated sources, we primarily use the eccentric version of the time-domain TEOBResumS model (called TEOBResumS-DALI) that can incorporate high eccentricities in the presence of aligned-spins (Chiaramello and Nagar, 2020; Nagar et al., 2021). Here we also use it for generating templates for BBH systems. Currently, neither of these waveform models incorporate the mean anomaly parameter and thus, all our injections and the results are obtained at an arbitrary, fixed zero mean anomaly. The waveforms are computed at a reference frequency of 20 Hz. Our key observations and insights are summarized as follows.
@@
To explore a given search parameter space, the PSO algorithm iteratively computes the templates points on the fly while optimizing a given detection statistic. This approach is different from that of a template-bank which depends on a fixed, pre-determined set of template points. The PSO algorithm thus, avoids the minimal-match criterion implemented in the template-bank searches but spends more time in exploring the region of the parameter space where the signal is more likely to be found. Some parts of the search parameter space generally require larger sampling than others. For example, the signal recovery in the low mass BNS region need more template points than that in the heavier BBHs. When data from more than one detector are available, the PSO algorithm can optimize a statistic computed over the detector network.
@@
Matched-filtering of the data is preceded by template waveform generation in the PSO framework. Eccentric templates are in general costlier than the quasi-circular ones impacting the overall cost of the search. We find that the above strategy is effective in recovering the effects of eccentricity in the presence of the aligned-spins. However, we are currently limited by the parameter space coverage of the waveform models and the speed of template generation on-the-fly.
@@
Eccentricity primarily modulates the inspiral component of the signals and the SNR gain is thus, mostly from the low frequency content of the signals. On the other end, truncating signals till Hz by downsampling reduces the cost of matched-filtering with little to no impact on recovered SNRs. If eccentricity is absent in the signals, the eccentric search approximately behaves like the usual aligned-spin search in terms of signal recovery.
@@ II.1 Detection principle
In this section, we briefly review the overall method for coincident detection of compact binary coalescences. The data recorded by a gravitational wave detector is a timeseries of discretely sampled strain, containing occasional gravitational signals in the detector’s noise. The detection problem is to essentially identify instances of transient astrophysical signals in the timeseries data. This is achieved by cross-correlating the strain data with a modeled form of the astrophysical signal, called a template waveform, . A template waveform depends on several parameters, modelling the coalescing binary. The cross-correlation is performed in the frequency domain where the noise power spectral density (PSD), is used to down-weight the effect of noise, in a process called matched-filtering. The output of matched-filtering is the signal-to-noise ratio, defined as,
where a tilde represents a function in the frequency domain. Whenever there is a close match between the true signal in the data and a template waveform, the SNR timeseries attains a peak at the arrival time of the signal in the detector as shown in Fig. 1. However at times, detector noise can also mimic GW signals by causing sharp peaks in the SNR timeseries. These occur mainly due to the non-stationarity and non-Gaussian nature of the detector noise and are called glitches. Glitches increase the chances of false alarms and affect the sensitivity to true astrophysical signals. The peaks in the SNR timeseries above a chosen threshold which arise due to the noise glitches and any potential astrophysical signals are considered as triggers. The triggers undergo several consistency checks to veto out possible glitches. To get rid of some of the obvious loud noise transients, the signal-based power discriminator is used (Allen, 2005). Other vetoing techniques are also available for the noise triggers that dodge the standard consistency checks, especially from short duration templates (Nitz, 2018). The discriminators are designed to down-rank noise triggers in a quantity called the reweighted SNR while preserving the original SNR values for the triggers with astrophysical origin (Usman et al., 2016). Remaining triggers are tested for time-coincidence with multiple detectors - the arrival times of astrophysical signals at any pair of detectors cannot differ by more than the time-of-flight for light between the detectors with some tolerance for the measurement errors. Events that qualify these tests are assigned a statistical significance based on the rate of false alarms, which is described in section II.4. The general detection problem is extensively discussed in the literature, for example (Sathyaprakash and Schutz, 2009; Schutz, 1997; Sathyaprakash and Dhurandhar, 1991).
@@ II.2 Swarm-intelligence
Identifying a potential signal in a given segment of data hinges on finding a close match between the signal and a template waveform. As described in the previous section, one needs to optimize the function as given in the Eq. (1) over the desired range of parameters . In this section, we describe how the proposed algorithm locates the optimal template position which is represented by . The algorithm begins with a fixed number of randomly distributed template points which we refer to as a swarm. The individual members of the swarm, which are known as particles, communicate and share the knowledge gained during the search process. Each particle with a position and a velocity , where is the particle identifier and is the dimension index, constitutes a candidate for the optimal template . The initial random distribution of the particles can be chosen to be uniform in each of the search dimensions. The role of the algorithm is to iteratively improve a given fitness function by proposing a new set of positions to the particles and computing the fitness function at the newly proposed locations. The swarm eventually converges over few tens of iterations (typically less than 20) while the value of the fitness function saturates at an optimal value. Note that the fitness function must be carefully chosen since its repeated evaluation can be costly. For example, the simplest choice is the matched-filter SNR maximized over time for a single detector with Gaussian noise. However, a noise glitch can cause a very large SNR value and can deceive the algorithm. Thus we replace the SNR with the reweighted SNR for the network of detectors for maximization, as discussed in Section-II.3. To take any subsequent step, a particle considers the following three factors-
@@
These three tendencies are reflected through the main three terms on the right-hand side of Eq. 2. So after computing the fitnesses of the individual particles at any given step, the algorithm identifies two quantities: (a) the template point with the largest fitness function among all the particles, which is called global-best or g-best, denoted as , and (b) the template point with the highest fitness achieved by the particle till the current step, which is called personal-best or p-best, denoted as . The overall velocity at step can be expressed as:
where and are uniformly distributed random numbers between and . The constant weights and control the degree of convergence and exploration of the swarm. A PSO search is usually stable within a broad range of values for these hyperparameters. A rough set of values is (, , ) = (0.7, 0.35, 0.4), but fine-tuning is possible depending upon the search. The positions of the particles are updated using the simple rule given below,
As an initial test of robustness of the PSO algorithm, we demonstrate the recovery of simulated BNS sources from a small population of about 500 injections in Gaussian noise. These are non-eccentric BNS injections with moderate aligned-spin components; the component masses are randomly drawn from . To locate the optimal template parameters for the HL detector network, we used 5k particles and allowed up to 15 steps, aggregated to about 75k template evaluations and that many matched-filter computations per detector per injection. The initial distribution of the particles are uniform for the search dimensions given by
This corresponds to an aligned-spin search using PSO. A summary of the simulation results are shown in Fig. 2. The injected coincident SNR is the coincident SNR obtained from the template with the injected parameters. Here an average loss of about 0.4% is observed in the recovered coincident SNR from that of the injected values. We also observed that to tackle less dense regions of the search space, like that of the NSBH and BBHs, requires even lesser sampling of the parameter space in general. The final g-best location in the parameter space represents an optimal solution with the obtained template parameters . Here the point estimates of the parameters describing the simulated sources are reasonably unbiased, notably the chirp mass, and the effective spin parameter, as shown in Fig. 2. The modifications to this algorithm to adapt to the real instrumental noise are described in the next subsection.
@@ II.3 Multiple detector optimization
When searching in the coincident mode, the processes described earlier in section II.2 can, in principle, proceed independently and in parallel for all the available detectors. However, in such a case, the templates are also computed independently for each detector and the point of convergence can differ due to several reasons. When projected from arbirary directions in the sky, the antenna pattern functions of the detectors modulate the incoming signal differently, thus effectively shifting the expected optimal point. In addition, the noise contributions from the individual detectors can slightly shift the convergence point. Otherwise, several local peaks form an elongated band of almost degenerate templates with similar SNRs as shown in Fig. 3, which can confuse the convergence. But a true gravitational wave signal should have the times of arrival and the source parameters consistent across all the available detectors. To ensure the second criterion, the same set of template parameters should optimize a fitness function computed for the detector network. Besides, for a given template, a stretch of data can contain several noise triggers, however, such triggers are expected to be uncorrelated in time and appear at just one detector. A data stretch here represents the data recorded by the available detectors within a fixed GPS start and end times. Thus, to minimize spurious optimization of glitch triggers, a natural choice is to jointly optimize the reweighted SNRs across the network of detectors. The joint fitness function for a network of detectors is given by,
@@
where is the reweighted SNR of the trigger most likely arising from an astrophysical signal at the th detector. Template points that trigger to a loud glitch in one of the detectors are downranked by the veto processes and the swarm further inhibits convergence to such template points over the detector network. An additional advantage of the above procedure is that a single set of computation of the template waveforms serves the entire network of detectors. It is especially useful when generating the template waveforms is expensive as is generally the case in an eccentric search. A similar strategy is also employed to optimize the network SNR of the candidates found in the PyCBC Live search (Canton et al., 2021).
The steps involved in the optimization process is described as follows. These are also summarized in Fig. 4. To demonstrate the process, we assume that the coincident strain data from different detectors exist in the computer storage. The data blocks for each detector are divided into segments overlapping with their immediate neighbours in time to account for the corrupted ends of the SNR timeseries due to filtering artifacts. The duration of the data stretch is chosen in a way that an in-band inspiral signal lies completely within a data stretch. The PSO algorithm explores each data stretch one after another. Data segments are downsampled till 1024 Hz as discussed later in Section III. PSDs used in the matched-filters are computed from the segments using Welch’s method (Welch, 1967).
@@
To evaluate the fitness function for a given data stretch at any template point, the template waveform at is constructed and cross-correlated with the stretch of data. This results into an SNR timeseries per detector whose ends are discarded from the either side. Triggers from the remaining portion of the SNR timeseries above a chosen threshold, typically , are examined. We compute the reweighted SNR values at the trigger points using the power -test. Additionally, we compute the ratio of the reweighted SNR and the SNR for the triggers and use it to rank them for their astrophysical origin. Triggers with a ratio larger than (referred to as primary triggers) are retained. The highest ranked trigger, i.e. the one with the highest ratio, from each detector is used to update the fitness function associated with a given particle as given in (Eq. 5). Note that we do not consider triggers with a ratio less than . If no such trigger exists for a given template point for the data stretch, the corresponding particle proceeds for optimization with its previous value of the fitness function.
@@
Based on the updated values of the fitness function of each particle, the p-best and the g-best locations and fitnesses are computed as described in subsection II.2. Following this, Eq. 2-3 are used to assign a new set of positions to the particles for the next iteration. The algorithm also checks whether any particle has drifted beyond the desired search boundaries- a reflecting boundary puts back an outlier inside in a symmetric position while flipping its velocity 111While this is suitable in most cases, we find that an adsorbing, boundary is more effective for the eccentric dimension towards the recovery of non-eccentric sources, where instead of putting the outlier back to a symmetric location, we put it on the boundary.. The algorithm additionally checks for a terminal condition of a convergence below a small mean radius of the swarm or a preset timeout, whichever is earlier. The boundary conditions can improve the termination and also ensures no hang-up occurs due to non-availability of template parameters outside the region allowed by a waveform model. Note that these steps, namely evolving the position-velocity, applying the boundary conditions and the termination checks are computationally trivial. At this point, the current iteration is complete.
The algorithm repeats this procedure for few tens of iterations and returns the final g-best template point for the data stretch. We further compute the triggers from the optimized g-best template (or the detection template) in a similar fashion as described earlier and keep all the triggers that survive the ratio cut. These triggers from the detection template (referred to as secondary triggers) are used to form coincidences in a post-processing step as described in the next subsection.